OpenStudy (anonymous):

Write the equation of the line in Slope-Intercept form, Then graph the equation by using either the slope-intercept equation or by using x- and y-intercepts. The line contains points (4, 6) and (-1, 2).

OpenStudy (anonymous):

Find the slope: change in y/change in x (y2-y1/x2-x1) 2-6/-1-4 = -4/-5 = 4/5 Plug into point-slope form: y-k=m(x-h) m= slope line passes through point (h,k) y-6=4/5(x-2) y-6=4/5x-1.6 y=4/5x+4.4 y-intercept: 4.4 x-intercept: 0=4/5x+4.4 -4.4=4/5x x=-5.5 Graph the line using these points in addition to the points you are given.

OpenStudy (anonymous):

Sooo lost....